# Acquiring Better Load Estimates by Combining Anomaly and Change-point Detection in Power Grid Time-series Measurements

### v1

@@ 4 Discussion
The current analysis considers one year of measurements for one station at a time. This is in line with the labeling procedure used by the grid operator prior to this research, where the data was labeled by hand during the first month of the year. As load estimates are used in long-term decision making, generally the speed of acquisition is not that important. However, should the availability of more recent load estimates be prioritised, the binary segmentation algorithm, which only works on static data, will need to be replaced by an online change point detection algorithm.
The methods employed for the change point and anomaly detection are all highly interpretable and fairly easily comprehensible. In recent years, machine learning has been widely applied in high-performance applications on time series. Examples are LSTM’s for anomaly detection Lindemann et al. (2021), and Meta’s (formerly Facebook’s) Prophet Taylor and Letham (2018) for simultaneous anomaly and change point detection. While promising, these methods would not be readily usable on this application. These methods require far more data than the unsupervised methods used in this research, something which is unfeasible due to time of data acquisition. While these methods can be trained on the unlabeled time series, we still need labels for hyperparameter optimization. Using the model with the best predictions doesn’t guarantee the best model for event detection, as an overparametrized model might learn to predict anomalies. Furthermore it has been shown that in many cases simple solutions work better than complex ones for time series anomaly detection Wu and Keogh (2021).
@@ 5 Conclusion
By having compared different methods and ensembling strategies we find that a combination of the well-known binary segmentation algorithm for change point detection and the tried statistical process control method for anomaly detection works best. The best ensembling strategy is a sequential ensemble, where the anomaly detector is applied after first segmenting the time series based on the obtained change points. The proposed methods are highly interpretable, a distinct advantage when this methodology is used in critical infrastructure planning.

### v3

@@ 4 Discussion
The current analysis considers one year of measurements for one station at a time. This is in line with the labeling procedure used by the grid operator prior to this research, where the data was labeled by hand during the first month of the year. As load estimates are used in long-term decision making, generally the speed of acquisition is not that important. However, should the availability of more recent load estimates be prioritized, the binary segmentation algorithm, which only works on static data, will need to be replaced by an online change point detection algorithm.
The methods employed for the change point and anomaly detection are all highly interpretable and fairly easily comprehensible. In recent years, machine learning has been widely applied in high-performance applications on time series. Examples are LSTM’s for anomaly detection (Lindemann et al., 2021), and Meta’s (formerly Facebook’s) Prophet (Taylor and Letham, 2018) for simultaneous anomaly and change point detection. While promising, these methods would not be readily usable on this application. These methods require far more data than the unsupervised methods used in this research, something which is unfeasible due to time of data acquisition. While these methods can be trained on the unlabeled time series, we still need labels for hyperparameter optimization. Using the model with the best predictions does not guarantee the best model for event detection, as an overparametrized model might learn to predict anomalies. Furthermore it has been shown that in many cases simple solutions work better than complex ones for time series anomaly detection (Wu and Keogh, 2021).
@@
It should be noted that while figures like Figure 7 can be interpreted fairly easily, it is not immediately clear why the thresholds are chosen like they are. One could change the thresholds for this figure and achieve a better filtering. When interpreting these figures, one should always note that the thresholds are optimized to find the best F1.51.51.51.5 score over all stations, rather than the single figure shown.
@@ 5 Conclusion
By having compared different methods and ensembling strategies we find that a combination of the well-known binary segmentation algorithm for change point detection and the tried statistical process control method for anomaly detection works best. The best ensembling strategy is a sequential ensemble, where the anomaly detector is applied after first segmenting the time series based on the obtained change points. The proposed methods are highly interpretable, a distinct advantage when this methodology is used in critical infrastructure planning. This high interpretability is a direct result of each underlying model being interpretable. SPC and BS are both simple, yet effective mathematical models. The strategies used for optimization are similarly straightforward, and can be visualized. Depending on underlying needs or when business priorities change the chosen threshold(s) for the algorithms can be adjusted based on the easily translatable precision and recall measures using figures such as Figure 5.
@@ Appendix A Evaluated and Best Hyperparameters
Appendix B AUC-ROC performance of each method
To provide insight into the ranking of the anomalies, so without explicit thresholding, we provide an additional plot of the area under the curve of the receiver-operating characteristic (AUC-ROC) in Figure 10.
Figure 10: A bar plot of the results of each method per length category in terms of the area under the curve of the receiver-operating characteristic (AUC-ROC).